OpenStudy (ilovecake):

What is the value of the variable? 12a = 48 a =

OpenStudy (midhun.madhu1987):

divide both sides by 12 and u get a.

OpenStudy (mathstudent55):

Since the variable "a" is being multiplied by 12, you need to do the opposite operation to multiplying by 12. What is the opposite operation to multiplication?

OpenStudy (ilovecake):

4?

OpenStudy (midhun.madhu1987):

yes

OpenStudy (mathstudent55):

12a = 48 12a/12 = 48/12 a = 4 You are correct.

OpenStudy (ilovecake):

ok thanks